OpenStudy (anonymous):

Solve X 23-x -----= 7 3

OpenStudy (anonymous):

2

OpenStudy (dumbcow):

1) multiply by 3 on both sides 23-x = 21 2) subtract 23 on both sides -x = -2 3) Flip the signs x=2

OpenStudy (anonymous):

Sweet thanks =) Would this one be also be 2 ? x 1 - = - 8 4

OpenStudy (anonymous):

yes sweet

OpenStudy (anonymous):

Im sorry =/ This isnt as easy for me as it is for you >.> And thanks.

OpenStudy (anonymous):

Its actually old an old assignment given to me for exercise but some of the questions i dont get.

OpenStudy (anonymous):

Im a year 7 and its creepy asking someone how old they are.

OpenStudy (anonymous):

that u in the pic?

OpenStudy (anonymous):

No its a cartoon O.o

OpenStudy (anonymous):

oh pellet.i ws beginnig to fancy the gal :)

OpenStudy (anonymous):

lol it's creepy to ask someone how old they are!!

OpenStudy (anonymous):

im not seven year old u noob. im in year seven like grade seven >.>

OpenStudy (anonymous):

Lol i agree with AnwarA it is creepy asking a stranger that and its wrong to ask a girl that too. Him1618 ur silly for starting to like someone as dumb as me XD

OpenStudy (anonymous):

hahaha..no i ws beginning to fancy the girl in the pic...

OpenStudy (anonymous):

Yeh shes pretty huh..You shuld see my other drawings too =P

OpenStudy (anonymous):

sure.....

OpenStudy (anonymous):

\[\frac{23-x}{3}=7\]\[23-x=21\]\[23-21=x\]\[x=2\]

OpenStudy (anonymous):

Yehp thats how i got it thanks

OpenStudy (anonymous):

did u got ur previous problem? regarding slope...i've ansered it..

OpenStudy (anonymous):

Yup got it just having a bit of trouble with sin cos and tan =/ I kinda know how to use them but they didnt give me all the degrees =C Just obvious one thats 90 degrees how do i work out the other two ?

OpenStudy (anonymous):

post ur exact problem...

OpenStudy (anonymous):

I know a = 90 But i dont know how to work out b and c I i shuld just asume that they are both 45

OpenStudy (anonymous):

what is the answer of 8+6?

OpenStudy (anonymous):

14 ?

OpenStudy (anonymous):

No No!

OpenStudy (anonymous):

it depends on....

OpenStudy (anonymous):

I dont understand ? and what does that have to do with the problem i just posted ?

OpenStudy (anonymous):

Hmmmm... some times 14 sometimes 2! just look at your clock!